# The Leftmost particle of branching subordinators

### v1

@@ Appendix A Construction of branching subordinators
    The finiteness comes from the fact that, for any fixed $t$, $\sharp\mathscr{N}^{a}(t)<\infty$ almost surely (see the discussion after Definition A.1) and that $t\mapsto\sharp\mathscr{N}^{a}(t)$ is almost surely non-decreasing by construction of $Y^{a}$. We now prove that positions of particles are distinct. We have $\phi(\lambda)=\phi^{a}(\lambda)+\int_{\mathscr{Q}^{\star}}\left(1-e^{-\lambda x_{1}}\right)\Lambda^{a}(\mathrm{d}\bm{x})$ by (42) and (43). Since $\Lambda^{a}(\mathscr{Q}^{\star})<\infty$ (see a little before (43)) the second term is bounded as $\lambda$ goes to infinity, so we get from Assumption 1 that $\phi^{a}(\lambda)$ is unbounded. By (43), if we had $\Lambda^{a}_{1}((0,\infty))<\infty$, then $\phi^{a}(\lambda)$ would be bounded. We thus have $\Lambda^{a}_{1}((0,\infty))=\infty$. By [21, Thm 27.4], the law of the subordinator with Lévy measure $\Lambda^{a}_{1}$ is continuous at any fixed $t$. Using that $(\xi_{u})_{u\in\mathscr{U}}$ are i.i.d subordinators with Lévy measure $\Lambda^{a}_{1}$, the independence, from the above construction, of $(\xi_{u})_{u\in\mathscr{U}}$ and $(\mathfrak{b}^{a}_{u})_{u\in\mathscr{U}}$, and the definition of $(Y^{a}_{u}(\cdot))_{u\in\mathscr{U}}$ we get that the following occurs with probability one: at all rational times $t$ (resp. at all times $\mathfrak{b}^{a}_{v}<\infty$), the random variables in the collection $\{Y^{a}_{u}(t),u\in\mathscr{N}^{a}(t)\}$ (resp. $\{Y^{a}_{u}(\mathfrak{b}^{a}_{v}-),u\in\mathscr{N}^{a}(\mathfrak{b}^{a}_{v}-)\}$) take pairwise distinct values. ∎
@@ Appendix B Linear behavior and justification of (13) and (14)
Recall that $(Y(n))_{n\in\mathbb{N}}$ is a branching random walk on the non-negative real line such that $Y(0)=0$ and it is supercritical, see Remark 1.1. Let us first assume that (6) holds for some $\theta\geq 0$. It is well known since the famous works of Hammersley [16], Kingman [18] and Biggins [10] that, if $-\infty<\kappa(\theta_{0})<0$ for some $\theta_{0}>0$ (allowing $\kappa(0)=-\infty$), then, $\mathbb{P}$-almost surely
@@
where the non-negativity comes from $\kappa(\cdot)$ being non-decreasing. In fact, the above result holds on the set of non-extinction $\{\forall n\in\mathbb{N};\;\mathscr{N}_{n}\not=\varnothing\}$, but as we have already mentioned previously, the condition $\Lambda(\{\emptyset\})=0$ in (3) says there is no killing, that is, by construction, $\mathbb{P}(\forall n\in\mathbb{N};\;\mathscr{N}_{n}\not=\varnothing)=1$. Using that $t\mapsto\underline{Y}(t)$ is almost surely a non-decreasing function, we immediately deduce (14).
One can see by construction that if we only require (4), then for any $a>a_{0}(\Lambda)$ (see Section 1.2), $\underline{Y}(t)\leq\underline{Y}^{a}(t)$ and $Y^{a}$ is a branching subordinator with characteristics $(d,\Lambda^{a})$ such that $\Lambda^{a}$ satisfies (6) for all $\theta\geq 0$. Moreover, $\kappa_{a}(0)=\Lambda(\{\emptyset\})-\int_{\mathscr{Q}}(\sum_{k\geq 2}\mathds{1}_{\{x_{k}<a\}})\Lambda(\mathrm{d}\bm{x})=-\int_{\mathscr{Q}}(\sum_{k\geq 2}\mathds{1}_{\{x_{k}<a\}})\Lambda(\mathrm{d}\bm{x})<0$ for all $a>a_{0}(\Lambda)$. Since $\kappa_{a}(\cdot)$ is continuous, one can always find $\theta_{a,0}>0$ such that $-\infty<\kappa_{a}(\theta_{a,0})<0$. Hence, thanks to (14), we have $\limsup_{t\to\infty}\underline{Y}(t)/t\leq\sup_{\lambda\geq\theta_{a,0}}\kappa_{a}(\lambda)/\lambda=\sup_{\lambda>0}\kappa_{a}(\lambda)/\lambda$ for all $a>a_{0}(\Lambda)$, that is, by (11)
$\limsup_{t\to\infty}\frac{\underline{Y}(t)}{t}\leq d+\inf_{a>a_{0}(\Lambda)}\sup_{\lambda>0}\frac{1}{\lambda}\Big{(}\int_{\mathscr{Q}}\big{(}1-e^{-\lambda x_{1}}\big{)}\Lambda(\mathrm{d}\bm{x})-M_{a}(\lambda)\Big{)}\in[d,\infty),$
and this yields (13). Still from the construction of the branching subordinator in Appendix A, each particle in the system moves with a drift $d$ and has only positive jumps and branchings with non-negative displacements. Therefore all particles in the system at time $t$ have a position larger or equal to $dt$ so $\liminf_{t\rightarrow\infty}\underline{Y}(t)/t\geq d$. Hence
$d\leq\liminf_{t\to\infty}\frac{\underline{Y}(t)}{t}\leq\limsup_{t\to\infty}\frac{\underline{Y}(t)}{t}\leq d+\inf_{a>a_{0}(\Lambda)}\sup_{\lambda>0}\frac{1}{\lambda}\Big{(}\int_{\mathscr{Q}}\big{(}1-e^{-\lambda x_{1}}\big{)}\Lambda(\mathrm{d}\bm{x})-M_{a}(\lambda)\Big{)}\in[d,\infty).$
In particular, if $d>0$, then the growth of $t\mapsto\underline{Y}(t)$ is linear.
@@ Appendix C An equivalent formulation of Assumption 2
    Choosing $\lambda=c_{1}a^{\sigma}$ in (49) and using (47) we get the finiteness of the limsup in (17). Then, from $M_{a}(\lambda)=\int_{[0,a)}e^{-\mathfrak{z}\lambda}\mu(\mathrm{d}\mathfrak{z})$, we see that
@@
    We now assume that Assumption 2 holds true for some $\sigma>0$ and we prove (47) and (48). For any $a>0$ and $\lambda\geq 0$, we have $M_{a}(\lambda)\leq\sum_{n=0}^{\lfloor a\rfloor}e^{-\lambda n}\mu([n,n+1])$. Choosing $\lambda=c_{1}a^{\sigma}$ for some $c_{1}$ strictly larger than the limsup in (17), we get $M_{a}(c_{1}a^{\sigma})\leq\sum_{n=0}^{\lfloor a\rfloor}e^{-c_{1}a^{\sigma}n}\mu([n,n+1])\leq\sum_{n=0}^{\infty}e^{-c_{1}n^{1+\sigma}}\mu([n,n+1])$. This sum is finite by (17) so (47) follows. Then, (49) translates into $\log(M_{a}(\lambda))\geq-\lambda a+\log(\mu([0,a)))$. Choosing $\lambda=c_{2}a^{\sigma}$ for some $c_{2}$ strictly smaller than the liminf in (17), we get $\log(M_{a}(c_{2}a^{\sigma}))\geq-c_{2}a^{1+\sigma}+\log(\mu([0,a)))$. From this and (17), we get (48). ∎
In light of the equivalence from Lemma C.1, the double appearance of $\sigma$ in (48) is natural. Indeed, (47) implies the finiteness of the limsup in (17) which in turn implies, via (50), that for any function $f(\cdot):[0,\infty)\rightarrow[0,\infty)$ we have $\limsup_{a\rightarrow\infty}\log(M_{a}(f(a)))/a^{1+\sigma}<\infty$. Therefore, (48) only says that a counterpart of (47) is satisfied.
Acknowledgments. The authors are grateful to Zhan Shi and Bastien Mallein for interesting discussions and advises. The first author benefited during the preparation of this paper from the support of Hua Loo-Keng Center for Mathematical Sciences (AMSS, Chinese Academy of Sciences) and from the National Natural Science Foundation of China (No. 12288201).

### v2

@@ Appendix A Construction of branching subordinators
The finiteness comes from the fact that, for any fixed $t$, $\sharp\mathscr{N}^{a}(t)<\infty$ almost surely (see the discussion after Definition A.1). We now prove that positions of particles are distinct. We have $\phi(\lambda)=\phi^{a}(\lambda)+\int_{\mathscr{Q}^{\star}}\left(1-e^{-\lambda x_{1}}\right)\Lambda^{a}(\mathrm{d}\bm{x})$ by (45) and (46). Since $\Lambda^{a}(\mathscr{Q}^{\star})<\infty$ (see a little before (46)) the second term is bounded as $\lambda$ goes to infinity, so we get from Assumption 1 that $\phi^{a}(\lambda)$ is unbounded. By (46), if we had $\Lambda^{a}_{1}((0,\infty))<\infty$, then $\phi^{a}(\lambda)$ would be bounded. We thus have $\Lambda^{a}_{1}((0,\infty))=\infty$. By [25, Thm 27.4], the law of the subordinator with Lévy measure $\Lambda^{a}_{1}$ is continuous at any fixed $t$. Using that $(\xi_{u})_{u\in\mathscr{U}}$ are i.i.d subordinators with Lévy measure $\Lambda^{a}_{1}$, the independence of $(\xi_{u})_{u\in\mathscr{U}}$ and $(\mathscr{P}_{u})_{u\in\mathscr{U}}$, and the definition of $(Y^{a}_{u}(\cdot))_{u\in\mathscr{U}}$ we get that almost surely, at time $t$, the random variables in the collection $\{Y^{a}_{u}(t);u\in\mathscr{N}^{a}(t)\}$ (resp. $\{Y^{a}_{u}(t-);u\in\mathscr{N}^{a}(t-)\}$) take pairwise distinct values. The last claim follows from the first claim and Fubini’s theorem. Indeed, If we denote by $A(t)$ (resp. $\tilde{A}(t)$) the event where the collection $\{Y^{a}_{u}(t);u\in\mathscr{N}^{a}(t)\}$ (resp. $\{Y^{a}_{u}(t-);u\in\mathscr{N}^{a}(t-)\}$) fails to take pairwise distinct values, then $\mathbb{E}[\int_{0}^{\infty}\mathds{1}_{A(t)}\mathrm{d}t]=\int_{0}^{\infty}\mathbb{P}(A(t))\mathrm{d}t=\int_{0}^{\infty}0\mathrm{d}t=0$ (resp. $\mathbb{E}[\int_{0}^{\infty}\mathds{1}_{\tilde{A}(t)}\mathrm{d}t]=0$). ∎
We end this section by proving Lemma 4.2 from Section 4.2. The proof is an easy consequence of the following fact on Poisson point processes.
Fact A.8.
Let $\mathscr{I}$ be a subinterval of $0,\infty)$, $\mathscr{V}$ be a finite set equipped with its counting measure $\Sigma$ and $\mathscr{A}$ be a measured set equipped with a finite measure $\nu$. We consider a Poisson point process $\mathscr{P}$ on $\mathscr{I}\times\mathscr{A}\times\mathscr{V}$ with intensity measure $\mathrm{d}t\otimes\nu(\mathrm{d}a)\otimes\Sigma(\mathrm{d}v)$. Now let $S\subset\mathscr{I}$ have null Lebesgue measure and $f$ be a measurable function from $\mathscr{I}\setminus S$ to $\mathscr{V}$. We then define a point process $\mathscr{P}^{\prime}$ on $\mathscr{I}\times\mathscr{A}$ by $\mathscr{P}^{\prime}=\{(t,a);\ t\notin S,(t,a,f(t))\in\mathscr{P}\}$. Then $\mathscr{P}^{\prime}$ is a Poisson point process on $\mathscr{I}\times\mathscr{A}$ with intensity measure $\mathrm{d}t\otimes\nu(\mathrm{d}a)$.
Recall the elements introduced in Step 1 of the above construction of $Y^{a}$. For $u\in\mathscr{U}$, let us denote by $\mathscr{P}^{k}_{u}$ the restriction of the Poisson point process $\mathscr{P}_{u}$ to $0,2^{k}]$. Let $\mathscr{G}_{k}$ be the sigma field generated by $\{(\xi_{u})_{u\in\mathscr{U}},(\mathscr{P}^{k}_{u})_{u\in\mathscr{U}}\}$. Recall the notations $\mathscr{Q}^{\star}$ from (3) and $\mathscr{U}_{k}^{a}$, $S_{br}^{a}$, $u(t)$, $\bm{x}(t)$, $S_{low,k}$ from Section 4.2 and note that the random set $\mathscr{U}_{k}^{a}$ is measurable with respect to $\mathscr{G}_{k}$. By the construction from Step 1 we get that, on the event $\{\sharp\mathscr{U}_{k}^{a}\geq 1\}$, the conditional distribution of the point process $\{(t,\bm{x}(t),u(t));\ t\in[2^{k},2^{k+1})\cap S^{a}_{br},u(t)\in\mathscr{U}_{k}^{a}\}$ (where the conditioning is with respect to $\mathscr{G}_{k}$) is that of Poisson point process on $2^{k},2^{k+1})\times\mathscr{Q}^{\star}\times\mathscr{U}_{k}^{a}$ with intensity measure $\mathrm{d}t\otimes\Lambda^{a}(\cdot\cap\mathscr{Q}^{\star})\otimes\Sigma$, where $\Sigma$ denotes the counting measure on $\mathscr{U}_{k}^{a}$. Let $S$ denote the set of times $t\in[2^{k},2^{k+1})$ where the leftmost particle, among all particles in $\mathscr{U}_{k}^{a}$, is not unique, i.e. $S:=\{t\in[2^{k},2^{k+1});\ \sharp\{u\in\mathscr{U}_{k}^{a};Y^{a}_{u}(t-)=\min_{v\in\mathscr{U}_{k}^{a}}Y^{a}_{v}(t-)\}\geq 1\}$. We also define a function $f$ from $2^{k},2^{k+1})\setminus S$ to $\mathscr{U}$ that, at each time, assigns the index of the leftmost particle, among all particles in $\mathscr{U}_{k}^{a}$, at that time, i.e. $f(t):=u$ such that $Y^{a}_{u}(t-)=\min_{v\in\mathscr{U}_{k}^{a}}Y^{a}_{v}(t-)$ (therefore the range of $f$ almost surely lies in $\mathscr{U}_{k}^{a}$). Thanks to Assumption 4, the last term in the right-hand side of (48) is null almost surely. Therefore, the collection of trajectories $\{(Y^{a}_{u}(t))_{t\geq 2^{k}};u\in\mathscr{U}_{k}^{a}\}$ is measurable with respect to $\mathscr{G}_{k}$ and so is the random couple $(S,f)$. Moreover, Lemma A.7 tells us that, almost surely, $S$ has null Lebesgue measure. Therefore, applying Fact A.8, we get that, almost surely on $\{\sharp\mathscr{U}_{k}^{a}\geq 1\}$, the conditional distribution of the point process $\{(t,\bm{x}(t));\ t\in S_{low,k}\}$ (where the conditioning is with respect to $\mathscr{G}_{k}$) has the claimed distribution. Since the conditional distribution does not depend on the particular realization in $\{\sharp\mathscr{U}_{k}^{a}\geq 1\}$, the result follows. ∎
@@ Appendix B Linear behavior and justification of (17) and (18)
Recall that $(Y(n))_{n\geq 0}$ is a branching random walk on the non-negative real line such that $Y(0)=0$ and it is supercritical, see Remark 1.1. Let us first assume that (6) holds for some $\theta\geq 0$. It is well known since the famous works of Hammersley [18], Kingman [20] and Biggins [11] that, if $\infty<\kappa(\theta_{0})<0$ for some $\theta_{0}>0$ (allowing $\kappa(0)=$-$\infty$), then, $\mathbb{P}$-almost surely
@@
where the non-negativity comes from $\kappa(\cdot)$ being non-decreasing. In fact, the above result holds on the set of non-extinction $\{\forall n\geq 0,\;\mathscr{N}_{n}\not$$\varnothing\}$, but as we have already mentioned previously, the condition $\Lambda(\{\emptyset\})=0$ in (3) says there is no killing, that is, by construction, $\mathbb{P}(\forall n\geq 0,\;\mathscr{N}_{n}\not=\varnothing)=1$. Using that $t\mapsto\underline{Y}(t)$ is almost surely a non-decreasing function, we immediately deduce (18).
One can see by construction that if we only require (4), then for any $a>a_{0}(\Lambda)$ (see Section 1.2), $\underline{Y}(t)\leq\underline{Y}^{a}(t)$ and $Y^{a}$ is a branching subordinator with characteristics $(d,\Lambda^{a})$ such that $\Lambda^{a}$ satisfies (6) for all $\theta\geq 0$. Moreover, $\kappa_{a}(0)=\Lambda(\{\emptyset\})-\int_{\mathscr{Q}}(\sum_{k\geq 2}\mathds{1}_{\{x_{k}<a\}})\Lambda(\mathrm{d}\bm{x})=-\int_{\mathscr{Q}}(\sum_{k\geq 2}\mathds{1}_{\{x_{k}<a\}})\Lambda(\mathrm{d}\bm{x})<0$ for all $a>a_{0}(\Lambda)$. Since $\kappa_{a}(\cdot)$ is continuous, one can always find $\theta_{a,0}>0$ such that $\infty<\kappa_{a}(\theta_{a,0})<0$. Hence, thanks to (18), we have $\limsup_{t\to\infty}\underline{Y}(t)/t\leq\sup_{\lambda\geq\theta_{a,0}}\kappa_{a}(\lambda)/\lambda=\sup_{\lambda>0}\kappa_{a}(\lambda)/\lambda$ for all $a>a_{0}(\Lambda)$, that is, by (12)
$\limsup_{t\to\infty}\frac{\underline{Y}(t)}{t}\leq d+\inf_{a>a_{0}(\Lambda)}\sup_{\lambda>0}\frac{1}{\lambda}\Big(\int_{\mathscr{Q}}\big(1-e^{-\lambda x_{1}}\big)\Lambda(\mathrm{d}\bm{x})-M_{a}(\lambda)\Big)\in[d,\infty).$
Still from the construction of the branching subordinator in Appendix A, each particle in the system moves with a drift $d$ and has only positive jumps and branchings with non-negative displacements. Therefore all particles in the system at time $t$ have a position larger or equal to $dt$ so $\liminf_{t\rightarrow\infty}\underline{Y}(t)/t\geq d$. Hence
$d\leq\liminf_{t\to\infty}\frac{\underline{Y}(t)}{t}\leq\limsup_{t\to\infty}\frac{\underline{Y}(t)}{t}\leq d+\inf_{a>a_{0}(\Lambda)}\sup_{\lambda>0}\frac{1}{\lambda}\Big(\int_{\mathscr{Q}}\big(1-e^{-\lambda x_{1}}\big)\Lambda(\mathrm{d}\bm{x})-M_{a}(\lambda)\Big)\in[d,\infty),$
and this yields (17). In particular, if $d>0$, then the growth of $t\mapsto\underline{Y}(t)$ is linear.
@@ Appendix C An equivalent formulation of Assumption 2
Choosing $\lambda=c_{1}a^{\sigma}$ in (51) and using (49) we get the finiteness of the limsup in (21). Then, again by definition of $M_{a}(\lambda)$,
@@
We now assume that Assumption 2 holds true for some $\sigma>0$ and we prove (49) and (50). For any $a>0$ and $\lambda\geq 0$, we have $M_{a}(\lambda)\leq\sum_{n=0}^{\lfloor a\rfloor}e^{-\lambda n}\mu([n,n+1])$. Choosing $\lambda=c_{1}a^{\sigma}$ for some $c_{1}$ strictly larger than the limsup in (21), we get $M_{a}(c_{1}a^{\sigma})\leq\sum_{n=0}^{\lfloor a\rfloor}e^{-c_{1}a^{\sigma}n}\mu([n,n+1])\leq\sum_{n=0}^{\infty}e^{-c_{1}n^{1+\sigma}}\mu([n,n+1])$. This sum is finite by (21) so (49) follows. Then, (51) translates into $\log(M_{a}(\lambda))\geq-\lambda a+\log(\mu([0,a)))$. Choosing $\lambda=c_{2}a^{\sigma}$ for some $c_{2}$ strictly smaller than the liminf in (21), we get $\log(M_{a}(c_{2}a^{\sigma}))\geq-c_{2}a^{1+\sigma}+\log(\mu([0,a)))$. From this and (21), we get (50). ∎ In light of the equivalence from Lemma C.1, the double appearance of $\sigma$ in (50) is natural. Indeed, (49) implies the finiteness of the limsup in (21) which in turn implies, via (52), that for any function $f(\cdot):[0,\infty)\rightarrow[0,\infty)$ we have $\limsup_{a\rightarrow\infty}\log(M_{a}(f(a)))/a^{1+\sigma}<\infty$. Therefore, (50) only says that a counterpart of (49) is satisfied.
Acknowledgments. The authors are grateful to Zhan Shi and Bastien Mallein for interesting discussions, advises, and careful reading. This work is supported by Beijing Natural Science Foundation, project number IS24067. The first author benefited during the preparation of this paper from the support of Hua Loo-Keng Center for Mathematical Sciences (AMSS, Chinese Academy of Sciences) and from the National Natural Science Foundation of China (No. 12288201).